\begin{thebibliography}{12}
%\bibitem{ABD1} A. Andrada,  M. L. Barberis and  I.  Dotti, Classification of abelian complex structures on 6-dimensional Lie algebras, {\em J. Lond. Math. Soc.} (2) {\bf 83} (2011), no. 1, 232--255.
%\bibitem{ABD} A. Andrada, M. L.  Barberis and I. Dotti, Abelian Hermitian Geometry, {\em Differential Geom. Appl.} {\bf 30} (2012),  509--519.
%\bibitem{Aus}
%L. Auslander, An exposition of the structure of solvmanifolds. I. Algebraic theory,   {\em Bull. Amer. Math. Soc.} {\bf 79}  (1973), no. 2, 227--261.
%
%\bibitem{BDM} M. L. Barberis, I. G. Dotti and  R.J. Miatello,  On
%certain locally homogeneous Clifford manifolds, \emph{Ann. Glob. Anal. Geom.} {\bf 13}
%(1995), 513--518.
%%
%\bibitem{BR}
%O. Baues and J. Riesterer, Virtually abelian K\"ahler and projective groups.
%{\em Abh. Math. Semin. Univ. Hambg.} {\bf 81} (2011), no. 2, 191--213.

\bibitem{AB1}
L. Alessandrini and G. Bassanelli, Metric properties of manifolds bimeromorphic to compact
K\"ahler spaces. {\em J. Diff. Geom.} {\bf 37} (1993), 95--121.

\bibitem{AB2}
L. Alessandrini and G. Bassanelli, Modifications of compact balanced manifolds. {\em C. R. Acad.
Sci. Paris S\'er. I Math.} {\bf 320} (1995), 1517--1522.

\bibitem{AV} A. Andrada and  R. Villacampa,  Abelian balanced hermitian  structures on unimodular Lie algebras, to apper in {\em Transform. Groups}, {\tt arXiv: 1412.7092.}.

\bibitem{Ivanov}
B. Alexandrov and S. Ivanov, Vanishing theorems on Hermitian manifolds, {\em Differential Geom. Appl.} {\bf 14} (2001), no. 3,
251--265.

\bibitem{BG} C. Benson and C. S. Gordon, K\"ahler and symplectic structures on nilmanifolds, {\em Topology}  {\bf 27} (1988), 513--518.

\bibitem{Bismut}
J. M. Bismut, A local index theorem for non-K\"ahler manifolds, {\em Math. Ann.} {\bf 284} (1989), 681--699.

\bibitem{Chiose}
 I. Chiose, Obstructions to the existence of K\"ahler structures on compact complex manifolds. {\em Proc. Amer. Math. Soc.} {\bf 142} (2014), no. 10, 3561--3568. 
%\bibitem{BDeG} D.  Burde and W.  A. de Graaf, Classification of Novikov algebras, {\em Appl. Algebra Engrg. Comm. Comput.} {\bf 24} (2013), no. 1, 1--15.

%\bibitem{Cavalcanti}  G.  Cavalcanti, SKT geometry,  {\tt arXiv:1203.0493}.

%\bibitem{Chu} B. Y. Chu,
%Symplectic homogeneous spaces,
%{\em Trans. Amer. Math. Soc.} {\bf 197} (1974), 145--159.
%
%\bibitem{CFK}
%S. Console, A. Fino and H. Kasuya, Modification and cohomology of solvmanifolds,  {\tt arXiv:1301.6042}.
%
%\bibitem{Dek} K. Dekimpe,  Solvable Lie algebras, Lie groups and polynomial structures,  {\em Compositio Math.} {\bf 121} (2000), no. 2,  183--204.
%
%\bibitem{dek} K.  Dekimpe,
%Semi-simple splittings for solvable Lie groups and polynomial structures, {\em Forum Math.} {\bf 12} (2000), no. 1, 77--96.
%
%\bibitem{Dix}
%J. Dixmier, Cohomologie des algebres de Lie nilpotentes, {\em Acta Sci. Math. Szeged} {\bf 16 } (1955), 246--250.
%
%\bibitem{DLZ}
%T. Draghici, T.-J. Li and W. Zhang: On the $J$-anti-invariant cohomology of almost complex 4-manifolds,   {\em Q. J. Math.} {\bf 64} (2013), no. 1, 83--111.
%
%\bibitem{EF} N. Enrietti and A. Fino, Special Hermitian metrics and Lie groups,  {\em Differential Geom. Appl.} {\bf 29}  (2011), suppl. 1, 211--219.
%
\bibitem{EFV}  N. Enrietti, A.  Fino and  L. Vezzoni, Tamed symplectic forms and strong K\"ahler with torsion metrics,  {\em J. Symplectic Geom. } {\bf 10}  (2012), no. 2, 203--223


\bibitem{FG} A. Fino and G. Grantcharov, On some properties of the manifolds with skew-symmetric torsion and holonomy $SU(n)$ and $Sp(n)$, 
{\em Adv. Math.} {\bf 189} (2004), 439--450. 

\bibitem{FPS} A. Fino, M.  Parton and S.  Salamon, Families of strong KT structures in six dimensions, {\em Comment. Math. Helv.}  {\bf 79}  (2004), no. 2, 317--340.

\bibitem{FT}  A. Fino and A. Tomassini,  A survey on strong KT structures, {\em  Bull. Math. Soc. Sci. Math. Roumanie}  (N.S.) {\bf 52} (100) (2009), no. 2, 99--116.

\bibitem{Hasegawa} K. Hasegawa, Minimal models of nilmanifolds, {\em Proc. Amer. Math. Soc.}  {\bf 106} (1989), 65--71.

\bibitem{FV}
A. Fino and L. Vezzoni, Special Hermitian metrics on compact solvmanifolds. {\em J. Geom. Phys.} {\bf 91} (2015), 40--53. 

\bibitem{GGPoon}
D. Grantcharov, G. Grantcharov and Y.S. Poon: Calabi-Yau connections with torsion on toric bundles. {\em J. Differential
Geom.} {\bf 78} (2008), no. 1, 13--32.

%\bibitem{Hasegawa}  K. Hasegawa, A note on compact solvmanifolds with K\" ahler structures,
%{\em Osaka J. Math.}  {\bf 43}  (2006), no. 1, 131--135.
%

\bibitem{IP} S.  Ivanov,  G. Papadopoulos, Vanishing theorems on  $( l \vert k)$-strong K\"ahler manifolds with torsion,  {\em Adv. Math. } {\bf 237} (2013), 147--164. 

%
\bibitem{Kas} H. Kasuya, Minimal models, formality and hard Lefschetz properties of solvmanifolds with local systems, {\em  J. Differential Geom.}  {\bf 93} (2013), no 2, 269--298.


%\bibitem{KVai} H. Kasuya, Vaisman metrics on solvmanifolds and Oeljeklaus-Toma manifolds,  {\em Bull. London Math. Soc.} {\bf 45} (2013),  15--26.
%
%\bibitem{KH} H. Kasuya, Hodge symmetry and decomposition on non-K\"ahler solvmanifolds, {\em J. Geom. Phys.} {\bf 76} (2014), 61--65.
%
%\bibitem{LZ}
%T.-J. Li and W. Zhang, Comparing tamed and compatible symplectic cones and cohomological properties of almost complex manifolds,  {\em Comm. Anal. Geom.} {\bf 17} (2009), no. 4, 651--683.
%


%\bibitem{OT}K. Oeljeklaus and M. Toma,  Non-K\"ahler compact complex manifolds associated to number fields,  {\em Ann. Inst. Fourier (Grenoble)} {\bf 55} (2005), no. 1, 161--171.

%\bibitem{Ov} G. Ovando, Invariant pseudo-K\"ahler metrics in dimension four, {\em J.  Lie Theory} {\bf 16} (2006), 371--391.


%\bibitem{OV} A. L. Onishchik and E. B. Vinberg (Eds), Lie groups and Lie algebras I
%\hspace{-.1em}I\hspace{-.1em}I,  Encyclopaedia of Mathematical Sciences, 41. Springer-Verlag, Berlin, 1994.
%
%\bibitem{P}
%T. Peternell, Algebraicity criteria for compact complex manifolds, {\em Math. Ann.} {\bf 275}
%(1986), no. 4, 653--672.

%\bibitem{Sn} J. E. Snow, Invariant complex structures on four dimensional solvable real Lie groups, {\em Manuscripta Math.} {\bf 66} (1990),
%397--412.

\bibitem{M}  M. L. Michelsohn, On the existence of special metrics in complex geometry. {\em Acta Math.} {\bf 149} (1982), no. 3--4, 261--295.

\bibitem{ST} J. Streets and  G. Tian, A Parabolic flow of pluriclosed metrics,  {\em Int. Math. Res. Not. IMRN} {\bf 2010} (2010), 3101--3133.
\bibitem{ST1} J. Streets and G. Tian, Regularity results for the pluriclosed flow, {\em Geom. and Top.} {\bf 17} (2013) 2389--2429.

\bibitem{LFY}
J. Fu, J. Li and S.-T. Yau: Balanced metrics on non-K\"ahler Calabi-Yau threefolds. {\em J. Differential Geom.} {\bf 90} (2012),
81--129.

\bibitem{ugarte}
L. Ugarte, Hermitian structures on six dimensional nilmanifolds, {\em Transf. Groups} {\bf 12}
(2007), 175--202.


\bibitem{Verb}  M. Verbitsky, Rational curves and special metrics on twistor spaces. {\em Geom. Topol.} {\bf 18} (2014), no. 2, 897Ð909.

%\bibitem{Yam}
%T. Yamada, Ricci flatness of certain compact pseudo-K\"ahler solvmanifolds, {\em J. Geom. Phys.} {\bf 62} (2012), no. 5, 1338--1345.
%
%\bibitem{Z}
%W. Zhang, From Taubes currents to almost K\"ahler forms, {\em Math. Ann.} {\bf 356} (2013), no. 9, 969--978.
\end{thebibliography}

\documentclass[11pt,oneside,leqno]{amsart}
\usepackage{amsxtra}
\usepackage{amsopn}
\usepackage{color}
\usepackage{amsmath,amsthm,amssymb}
\usepackage{amscd}
\usepackage{amsfonts}
\usepackage{latexsym}
\usepackage{verbatim}

\definecolor{verde}{rgb}{0.5,.7,.2}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{definition}[theorem]{Definition}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{remark}[theorem]{Remark}
\newtheorem{example}[theorem]{Example}
\newtheorem{question}[theorem]{QUESTION}
\newtheorem{remark-question}[section]{Remark-Question}
\newtheorem{conjecture}[section]{Conjecture}

\newcommand\C{{\mathbb C}}

\newcommand\R{{\mathbb R}}

\newcommand{\ov}[1]{\overline{ #1}}

\title[On the existence of  balanced and SKT metrics]{On the existence of  balanced and SKT metrics on nilmanifolds}
\author{Anna Fino and Luigi Vezzoni}
\address{Dipartimento di Matematica G. Peano \\ Universit\`a di Torino\\
Via Carlo Alberto 10\\
10123 Torino\\ Italy}
\email{annamaria.fino@unito.it}
 \email{luigi.vezzoni@unito.it}

\subjclass[2010]{Primary 32J27; Secondary 53C55, 53C30}
\thanks{This work was partially supported by the project PRIN {\em Variet\`a reali e complesse: geometria, topologia e analisi armonica},  the project FIRB {\em Differential Geometry and Geometric functions theory} and by GNSAGA (Indam) of Italy.\\
Keywords and phrases: {\em Special Hermitian metrics, nilmanifolds}
}

\begin{document}

\begin{abstract} On a complex manifold  an Hermitian metric which is simultaneously SKT and balanced has to be necessarily K\"ahler.  It  has been conjectured that  if a  compact complex manifold  $(M, J)$ has an   SKT metric and    a  balanced metric both compatible with $J$, then $(M, J)$   is necessarily K\"ahler.  
We show that  the conjecture is true for  nilmanifolds.
\end{abstract}

\maketitle

\section{Introduction}

A  Riemannian metric $g$  on a complex manifold $(M,J)$ is {\em compatible} with $J$  (or {\em $J$-Hermitian})  if  $g (J \cdot, J\cdot ) = g (\cdot, \cdot)$.
In the present paper we focus on the existence of {\em special} Hermitian metrics on complex manifolds. More precisely,  we study the existence of {\em strong K\"ahler with torsion} (shortly SKT) and balanced metrics compatible with the same complex structure. 
We recall that a  $J$-Hermitian metric    is called {\em SKT} (or {\em pluriclosed}) if its fundamental form $\omega$ satisfies 
$$
\partial \bar \partial \omega=0\,,
$$
while $g$ is called {\em balanced} if $\omega$ is co-closed, i.e. 
$$
d^*\omega=0,
$$
where $d^*$ denotes the formal adjoint operator of $d$ with respect to the metric $g$.

SKT metrics were introduced by Bismut in \cite{Bismut} and further studied in many papers (see e.g. \cite{FV,GGPoon,FT,LFY,ST,ST1,Verb}  and the references therein), while 
balanced metrics were introduced and firstly studied by Michelsohn in \cite{M}, where their existence is characterized in terms of currents. In a subsequent paper Alessandrini and Bassanelli 
proved that modifications of compact balanced manifolds are always balanced (see \cite{AB1,AB2}) showing a powerful tool for finding examples of balanced manifolds. 

It is well-known that if an Hermitian metric $g$ is simultaneously SKT and balanced, then it is necessarily K\"ahler (see e.g. \cite{Ivanov}).   This result has been generalized in  \cite{IP}  showing that  a compact  SKT conformally balanced  manifold has to be K\"ahler. The next conjecture was stated in \cite{FV} and it is about the existence of an SKT metric and a balanced metric both compatible with the same complex structure:

\medskip
\noindent {\bf Conjecture.} {\em Every compact complex manifold admitting both an SKT metric and a balanced metric is necessarily K\"ahler.} 

\medskip
The conjecture has been implicitly already proved in literature in some special cases. For instance,
Verbitsky has showed  in \cite{Verb} that the twistor space of a compact, anti-self-dual Riemannian manifold has
no SKT metrics unless it is K\"ahlerian and Chiose has obtained in \cite{Chiose} a similar result for non-K\"ahler
complex manifolds belonging to the Fujiki class. Furthermore, Li, Fu and Yau have proved in \cite{LFY} that some new examples of SKT manifolds do not admit any balanced metric.
A natural  source of non-K\"ahler manifolds admitting balanced metrics and SKT metrics are given by  {\em nilmanifolds}, i.e.   by compact manifolds obtained as quotients
of a  simply connected nilpotent Lie group $G$ by a co-compact lattice $\Gamma$.   It is well known that a nilmanifold  cannot admit K\"ahler  structures unless it a torus (see for instance \cite{BG,Hasegawa}).
The aim of the paper is to show that the conjecture is true when $(M,J)$ is a {\em complex nilmanifold}.  By {\em complex nilmanifold} we refer to   a nilmanifold equipped with an {\em invariant} complex structure $J$,  i.e. endowed  with a complex structure  induced  by a left-invariant complex structure on $G$. 

Our result is the following        

\begin{theorem}\label{main}
Let $M=G\backslash\Gamma$ be a nilmanifold equipped with an invariant complex structure $J$.  Assume that $(M,J)$ admits a balanced metric $g$ and an SKT metric $g'$ both compatible with $J$. Then $(M,J)$  is a complex torus. 
\end{theorem}
The theorem is trivial in dimension $6$ in view of the main result in \cite{FPS} and it was already proved in \cite{FV} when the nilmanifold has dimension $8$ by using a classification result proven in \cite{EFV}. 

\section{Proof of Theorem \ref{main}}

In order to prove  Theorem \ref{main}   we need the following  lemmas
\begin{lemma}\label{2.1}
Let $(M=G/\Gamma,J)$ be a complex nilmanifold. 
\begin{itemize}
\item If $(M,J)$ has a balanced metric, then it has also an {\em invariant} balanced metric \cite{FG}. 

\item If $(M,J)$ has an SKT metric, then it has also an {\em invariant} SKT metric \cite{ugarte}.
\end{itemize}
\end{lemma}

\begin{lemma}[\cite{EFV}]\label{2.3}
Let $(M=G/\Gamma,J)$ be a complex nilmanifold of real  dimension $2n$. If $(M,J)$ has an SKT metric, then $G$ is (at most) $2$-step nilpotent, and there exists a complex $(1,0)$-coframe $\{\alpha^1,\dots,\alpha^n\}$ on $\mathfrak g$ satisfying the following structure equations$$
\left \{ \begin{array}{l}
d \alpha^j= 0, \quad j = 1, \ldots, k,\\
d \alpha^j  = \sum_{r, s =1}^k  \left(\frac12 c_{rs}^j  \alpha^{r}\wedge \alpha^s + c_{r \overline {s}}^j   \alpha^{r}\wedge \bar \alpha^{s}\right),\, j=k+1,\dots,n \,,
\end{array} \right.
$$
for some $k\in \{1,\dots,n-1\}$ and  with $c_{rs}^j, c_{r \overline {s}}^j \in \C$.
\end{lemma}

Now we can prove Theorem \ref{main}

\begin{proof}[Proof of Theorem $\ref{main}$]
Suppose that $(M=G/\Gamma, J)$ is not a complex torus, i.e. that $G$ is not abelian
and denote by $\mathfrak g$ the Lie algebra of $G$.  Assume that $(M,J)$ admits a balanced metric $g$ and also an SKT metric $g'$ both compatible with $J$. 
Then in view of Lemma \ref{2.1}, we may assume both $g$ and $g'$ invariant and regarding them as scalar products on $\mathfrak g$. This allows us to work at the level of the Lie algebra $\frak g$. 
As a consequence of Lemma \ref{2.3}, the existence of the {\rm SKT} metric implies that Lie algebra  $\frak g$ is $2$-step nilpotent and  that $(\frak g,J)$ has a $(1,0)$-coframe 
$\{\alpha^1, \ldots ,\alpha^n\}$ satisfying  the following structure equations 
\begin{equation}\label{streq}
\left \{ \begin{array}{l}
d \alpha^j= 0, \quad j = 1, \ldots, k,\\
d \alpha^j  = \sum_{r, s =1}^k  \left(\frac12 c_{rs}^j  \alpha^{r s} + c_{r \overline {s}}^j   \alpha^{r{\overline{s}}}\right).
\end{array} \right.
\end{equation}
for some $k\in \{1,\dots,n-1\}$ and with $c_{rs}^j, c_{r \overline s}^j \in \C$. Here we use the notation $\bar \alpha^i=\alpha^{\bar i}$ and $\alpha^{r_1\cdots r_p\bar s_1\cdots \bar s_q}=
\alpha^{s_1}\wedge \dots \alpha^{s_p}\wedge \alpha^{\bar s_1}\wedge \dots \alpha^{\bar s_q}$.
We may assume without restrictions that   the coframe $\{\alpha^i\}$ is unitary with respect to the balanced metric $g$. Indeed,  
since in the Gram-Schmidt process  the spaces spanned by the first $r$ elements of the original basis are preserved, we can modify the coframe $\{\alpha^r\}$ making it unitary with respect to $g$ and satisfying the same structure equations as in \eqref{streq} with different structure constants. 
In this way $g$ writes as 
$$
g=\sum_{r=1}^n\alpha^r\otimes \alpha^{\bar r}
$$
and the balanced condition can be written in terms of $c_{r\bar s}^l$'s as 
\begin{equation}\label{balanced}
\sum_{r=1}^k c_{r \ov r}^ l =0,
\end{equation}
for every $l > k$ (see also \cite[Lemma 2.1]{AV}). 

Next we focus on the SKT metric $g'$. Since  $\mathfrak g$ is $2$-step nilpotent, we have 
 $\partial\bar \partial \alpha^r=0$, $r=1,\dots,n$.
If $$
\omega'=\sum_{i,j=1}^n a_{i\bar j} \alpha^i\wedge \alpha^{\bar j}
$$
is the fundamental form of $g'$, then the SKT condition $\partial \bar \partial \omega'=0$ writes as 
\begin{equation}\label{SKT}
\sum_{i,j=k+1}^n    a_{i \overline j}  ( \overline \partial \alpha^i \wedge   \partial  \alpha^{\ov j}  - \partial \alpha^i \wedge \overline  \partial \alpha^{\ov j}) =0\,. 
\end{equation}
Equation \eqref{SKT} can be written in terms of the  structure constants as 
$$
\sum_{i,j=k+1}^n \sum_{r, s,u,v =1}^k   a_{i \overline j} \left(c_{r\bar v}^i \bar c_{u\bar s}^{ j}+\frac14c_{rs}^i\bar c_{uv}^{j}\right)   \alpha^{rs \bar u\bar v}  =0.
$$
By  considering the component along $ \alpha^{rs \bar s\bar r}$ in the above expression we get 
$$
\sum_{i,j=k+1}^n  a_{i \overline j}\left( c_{r\bar r}^i \bar c_{s\bar s}^{ j}+\frac14c_{rs}^i\bar c_{sr}^{j}
-c_{s\bar r}^i \bar c_{s\bar r}^{ j}-\frac14c_{sr}^i\bar c_{sr}^{j}
+c_{s\bar s}^i \bar c_{r\bar r}^{ j}+\frac14c_{sr}^i\bar c_{rs}^{j}
-c_{r\bar s}^i \bar c_{r\bar s}^{ j}-\frac14c_{rs}^i\bar c_{rs}^{j}\right)=0\,,
$$
i.e. the condition
\begin{equation} \label{coefficient}
\sum_{i,j=k+1}^n  a_{i \overline j}\left(c_{r\bar r}^i \bar c_{s\bar s}^{ j}
-c_{s\bar r}^i \bar c_{s\bar r}^{ j}+c_{s\bar s}^i \bar c_{r\bar r}^{ j}
-c_{r\bar s}^i \bar c_{r\bar s}^{ j}+c_{rs}^i\bar c_{sr}^{j}\right)=0\,.
\end{equation}
Now by taking the sum of  \eqref{coefficient} for $r,s=1,\dots k$ and keeping in mind the balanced assumption \eqref{balanced} we get 
\begin{equation}\label{SKTnew}
\sum_{i,j=k+1}^n \sum_{r,s=1}^k  a_{i \overline j}\left(
2c_{s\bar r}^i \bar c_{s\bar r}^{ j}
+c_{rs}^i\bar c_{rs}^{j}\right)=0\,.
\end{equation}
Let us consider now the $(1,0)$-vectors $X_{sr}$ and $X_{r\bar s}$ on $\mathfrak g$ defined by 
$$
X_{rs}=\sum_{i=k+1}^nc_{rs}^iX_i\,,\quad X_{r\bar s}=\sum_{i=k+1}^n  \sqrt{2}\,  c_{r\bar s}^iX_i
$$
where $\{X_1,\dots,X_n\}$ is the dual frame to  $\{\alpha^1,\dots,\alpha^n\}$. 
We have 
$$
\sum_{r,s=1}^k\left(\omega'(X_{rs},\bar X_{rs})+ \omega'(X_{r\bar s},\bar X_{r\bar s})\right)=\sum_{i,j=k+1}^n \sum_{r,s=1}^k  a_{i \overline j}\left(
2c_{s\bar r}^i \bar c_{s\bar r}^{ j}
+c_{rs}^i\bar c_{rs}^{j}\right)
$$
and so equation \eqref{SKTnew} implies $X_{rs}=X_{r\bar s}=0$ for every $r,s=1,\dots, k$. So every $\alpha^r$ is closed and $\mathfrak g $ is abelian, contradicting the first assumption in the proof. 
\end{proof}

\begin{remark}
{\em Theorem \ref{main} is trivial in dimension $6$. Indeed, in view of \cite{FPS} and Lemma \ref{2.1}, if a $6$-dimensional complex nilmanifold  $(M, J)$ has an SKT metric, then every  invariant   $J$-Hermitian  metric on $(M,J)$  is SKT and so every  invariant  balanced metric compatible with $J$ is automatically K\"ahler.  More generally, the same argument works when $M$ has arbitrary  real dimension $2n$, but $k=n-1$. }
\end{remark}

\begin{remark}{\em 
Another case where Theorem \ref{main} is trivial is when $k=1$. In this case  the Lie algebra  $\frak g$ is necessary isomorphic to $\frak h_3^{\R} \oplus \R^{2n - 3}$, where  $\frak h_3^{\R}$ is the real $3$-dimensional Heisenberg algebra,  and the result follows.
}

\end{remark}

\noindent {\it Acknowledgements}. We would like to thank Luis Ugarte and Raquel Villacampa for useful comments on the paper.

\end{document}